Density of dry air containing only N2 and O2 is 1.15g/L at 740 mm of Hg and 300k. What is ℅ composition of N2 by weight in the air.
1
Expert's answer
2019-07-01T03:55:09-0400

Solution.

"p \times V = \frac{m}{M} \times R \times T"

"\rho = \frac{m}{V}"

"p = \frac{\rho \times R \times T}{M}"

"M = \frac{\rho \times R \times T}{p}"




"M = \frac{g \times J \times K}{Liters \times mole \times K \times Pa}"




"Pa = \frac{J}{m^3}""m^3 = 1000 \times Liters"


"M = \frac{g \times J \times 1000 \times Liters}{Liters \times mole \times J}"

M = 29.06 g/mole


"w(N2) = \frac{28.01}{29.06}"

w(N2) = 96.39 %

Answer:

w(N2) = 96.39 %